# Topology of Discrete Quantum Feedback Control

### v1

@@ I Introduction
Topology plays a key role in classification and characterization of diverse physical phenomena that are robust against local disturbances. In particular, topological phases of matter have been discussed in a wide variety of quantum and classical systems, which are characterized by global geometric structures of eigenstates rather than local order parameters [1, 2, 3, 4, 5, 6]. A quintessential example of topological quantum phenomena is the quantum Hall effect, where chiral transport immune to disorder emerges along the edges of a system [7, 8, 9, 10, 11]. Furthermore, topological phases are enriched by symmetry. For example, time-reversal symmetry permits helical edge states, where time-reversal partners flow in opposite directions with no net mass current which is known as the quantum spin Hall effect [12, 13, 14, 15, 16]. Classification of phases of matter through integration of topology and symmetry has been of central interest in condensed matter physics [17, 18, 19].
@@
The general concept of topology can also be utilized to classify nonequilibrium dynamics. Examples include periodically driven (Floquet) systems characterized by topology of unitary time-evolution operators [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31] and topological phases of open systems described by non-Hermitian Hamiltonians [32, 33, 34, 35, 36, 37, 38]. Notably, out-of-equilibrium driving not only provides a useful tool for controlling topological phases, but also induces genuinely dynamical topological phases that has no static counterpart [22, 39, 28, 33]. Finding a new control method for topological phases thus expands the scope of research of topological phases of matter.
@@
Feedback control provides a versatile tool for a wide range of applications in science and technology [40]. In particular, in the research arena of high-precision measurement and control of quantum systems, feedback control has achieved suppression of quantum noise [41], stabilization of desired states [42, 43], cooling [44, 45, 46], and quantum error correction [47, 48, 49, 50]. Feedback control also allows operations that would otherwise be impossible. An illustrative example is Maxwell’s demon, which was originally conceived as an intelligent being that can decrease entropy of a system against the second law of thermodynamics [51]. In modern language, Maxwell’s demon can be formulated as a feedback controller [52, 53], and various types of Maxwell’s demon have experimentally been realized in both classical [54, 55, 56, 57] and quantum [58, 59, 60, 61, 62, 63] systems. In view of versatile applicability of feedback control, we are naturally led to the following question: can we utilize feedback control to create and manipulate nonequilibrium topological phases? This is the problem we wish to address in this paper.
@@
In this paper, we present a class of dynamical topological phases that can be realized with feedback control of quantum systems. Here, a unique role played by feedback control is that one can choose operations in such a manner that they depend on measurement outcomes. This flexibility enables versatile real-time control of topological phases that would be hard to achieve by unitary dynamics in which the driving protocol is predetermined and cannot depend on the state of a system under control. Furthermore, feedback control enables active control of a quantum system, in contrast with passive control of open quantum systems by engineered dissipation.
@@
Figure 1: (a) Schematic illustration of the eigenspectrum (shown by a blue curve and two blue dots) of a quantum channel for topological quantum feedback control. The spectrum involves a closed loop characterized by a winding number and highly degenerate zero eigenvalues due to local measurement through which nonlocal coherence vanishes. The unit eigenvalue corresponds to a steady state. (b) Discrete quantum feedback control is comprised of a quantum measurement and a subsequent operation depending on the measurement outcome. Topologically nontrivial winding of the eigenspectrum of a quantum channel [see (a)] implies chiral (unidirectional) transport of a particle on a lattice due to feedback control.
@@
Table 1: Comparison of different classes topological phases. The first row shows five classes of topological phases characterized by the topology of relevant operators which are shown in the second row. The third row indicates whether the dynamics is discrete or continuous. The fourth row shows whether the system is closed or open. The fifth row shows whether a quantum state is pure or mixed.
@@
Quantum feedback control creates a class of nonequilibrium open quantum systems distinct from those described by non-Hermitian Hamiltonians [33, 34, 35] and Lindblad master equations [64, 65, 66, 67]. We summarize the comparison of different classes of topological phases in Table 1. In particular, discrete quantum feedback control concerns discrete time evolution from an initial quantum state to a post-feedback quantum state. Such time evolution cannot in general be expressed by using a generator of continuous time evolution. In fact, as shown in Sec. III, CPTP maps with local Kraus operators have highly degenerate zero eigenvalue, and therefore cannot be described by using any generator [see Eq. (39)]. Physically, the existence of zero eigenvalues can be understood as a consequence of local measurement through which nonlocal coherence in an extended system vanishes [see Fig. 1(a) and Sec. III]. A non-Hermitian Hamiltonian can only describe the dynamics of pure states and thus requires postselection of particular measurement outcomes [68]. The Lindblad quantum master equation describes the dynamics of mixed states but does not refer to measurement outcomes [40]. In contrast, nonequilibrium topological phases created by quantum feedback control are realized by a nontrivial interplay between quantum measurement and feedback operations. We note that non-Hermitian topology has been applied to quantum channels in Ref. [33]. However, its application beyond zero-dimensional systems is nontrivial; in fact, the locality of Kraus operators is the key for defining topological invariants, as discussed in Sec. III.
@@
We present prototypical models of topological feedback control by using quantum Maxwell’s demons, which we shall refer to as topological Maxwell’s demons. In a one-dimensional case, a topological Maxwell’s demon causes feedback-assisted chiral transport due to topology. We find that the topological Maxwell’s demon can generate topologically distinct phases by changing the duration of feedback control. We also show that feedback control can induce the non-Hermitian skin effect, which is a hallmark of dynamical topology unique to open systems [69, 70, 33, 71, 72, 73, 38, 74]. The chiral transport caused by the topological Maxwell’s demon is protected by topology and hence robust against noise and decoherence during the feedback process.
@@
Symmetry plays a pivotal role in classification of topological phases [18, 17, 19, 27, 28, 33, 34, 35, 65, 37, 66, 67]. Here, we perform a symmetry classification of quantum feedback control. We find that only a subset of the 38 symmetry classes of non-Hermitian matrices [75, 34, 35] is compatible with quantum channels with projective measurements. We identify ten relevant symmetry classes of quantum channels for discrete quantum feedback control (see Table 2), and provide examples of feedback control for each class. In particular, we propose a quantum Maxwell’s demon with a two-point measurement as an example of symmetry-protected topological feedback control. This demon generates helical spin transport protected by symmetry, which may be regarded as a counterpart of the quantum spin Hall effect in feedback-controlled quantum systems.
@@
Table 2: Ten-fold symmetry classification of discrete quantum feedback control with projective measurement. Each symmetry class is defined according to the presence or absence of $C,K$, and $Q$ symmetries in the CPTP map. Here, $C,K$, and $Q$ symmetries of a CPTP map $\mathcal{E}$ are defined by $\mathcal{V}_{C}\mathcal{E}^{T}\mathcal{V}_{C}^{-1}=\mathcal{E},\mathcal{V}_{K}\mathcal{E}^{*}\mathcal{V}_{K}^{-1}=\mathcal{E}$, and $\mathcal{V}_{Q}\mathcal{E}^{\dagger}\mathcal{V}_{Q}^{-1}=\mathcal{E}$, respectively, where $\mathcal{V}_{X}$ ($X=C,K,Q$) is a unitary superoperator. In the Table, 00 indicates the absence of symmetry, and ±$\pm 1$ indicates the presence of symmetry with $\mathcal{V}_{X}\mathcal{V}_{X}^{*}=\pm\mathcal{I}$, where $\mathcal{I}$ denotes the identity superoperator. The labeling of the symmetry classes follows Ref. [34], where psH denotes pseudo-Hermiticity and the subscript $+$ (--) of psH indicates that the unitary superoperator for $C$ or $K$ symmetry commutes (anticommutes) with that of the pseudo-Hermiticity. In addition to the symmetry classes of a CPTP map $\mathcal{E}$, the corresponding ${}^{\dagger}$ classes of $i\mathcal{E}$ are shown in the parentheses. In the fifth column, examples of conditions for feedback (FB) control in each symmetry class are shown (see Sec. VI for detailed discussions). CPTP maps with $C$ or $Q$ symmetry can be realized with feedback control with a two-point measurement (2PM). Symmetry classes without $K$ symmetry with $\mathcal{V}_{K}\mathcal{V}_{K}^{*}=+\mathcal{I}$ require unitary symmetry (US) due to the intrinsic modular conjugation symmetry of CPTP maps. ${}_{-}$ denotes antiunitary symmetry whose square gives $-1$. Some symmetry classes may be realized with an additional feedback control that enforces the symmetry condition (see examples in Sec. VII). The rightmost column shows concrete models of topological feedback control, which are presented in Sec. IV and Sec. VII.
@@
${}^{\dagger}$ $+1$ 00 00 US, 2PM, and reciprocity
${}_{+}$ (${}^{\dagger}$) $+1$ $+1$ $1$ 2PM and reciprocity
AI (${}^{\dagger}$) 00 $+1$ 00 No symmetry or particle-hole symmetry Chiral Maxwell’s demon / SSH demon
${}_{-}$ (${}^{\dagger}$) $-1$ $+1$ $1$ 2PM, reciprocity, and additional FB Helical Maxwell’s demon
${}^{\dagger}$ $-1$ 00 00 US, 2PM, reciprocity, and additional FB $\mathbb{Z}_{2}$ demon
${}_{+}$ (${}^{\dagger})$ $-1$ $-1$ $1$ US, 2PM, ${}_{-}$, reciprocity, and additional FB
AII (${}^{\dagger}$) 00 $-1$ 00 US and ${}_{-}$
${}_{-}$ (${}^{\dagger}$) $+1$ $-1$ $1$ US, 2PM, ${}_{-}$, and reciprocity
The rest of this paper is organized as follows. In Sec. II, we introduce quantum channels that describe discrete quantum feedback control. In Sec. III, we use the Bloch theory of translationally invariant quantum channels to show that topological invariants of quantum channels can be defined when Kraus operators are local in space. In Sec. IV, we construct an example of topological feedback control, which shows feedback-induced chiral transport due to its point-gap topology. In Sec. V, we examine symmetry of discrete quantum feedback control and identify ten symmetry classes that are compatible with projective measurement. We also discuss how to construct feedback control in each symmetry class in Sec. VI. On the basis of the symmetry classification, we provide various models of symmetry-protected topological feedback control in Sec. VII. In particular, we present topological feedback control protected by transposition symmetry which exhibits helical spin transport due to an interplay between symmetry and topology. In Sec. VIII, we extend our theory to the case of nonunitary feedback control, thereby showing robustness of topological feedback control against dissipation. We also propose feedback control with engineered dissipation, where dissipation is used as an essential tool to achieve feedback control. We conclude this paper in Sec. IX with an outlook. In Appendix A, we present the detail of the Bloch theory of quantum channels. In Appendix B, we describe the Bloch theory of quantum feedback control with projective measurement. In Appendix C, we discuss a spectral property of quantum channels for projective measurement. In Appendix D, we summarize the correspondence between symmetry of time-evolution operators and that of their generators. In Appendix E, we present a supplementary discussion of symmetry-preserving boundary conditions for a $\mathbb{Z}_{2}$ Maxwell’s demon.
@@ II Discrete quantum feedback control
We consider discrete quantum feedback control in which a feedback cycle consists of quantum measurements and subsequent unitary operations that are conditioned on measurement outcomes. Let $\hat{\rho}_{\mathrm{i}}$ be the initial density matrix of a system. We first perform a quantum measurement described by a set of measurement operators $\{\hat{M}_{m}\}$ subject to
@@
where $\hat{H}_{m}(t)$ is the feedback Hamiltonian which depends on the measurement outcome $m$ and $\mathcal{T}$ is the time-ordering operator. We set $\hbar=1$ throughout this paper. If we take an ensemble average over measurement outcomes, the final density matrix after one feedback loop is given by [52]
@@
If a feedback cycle is repeated, the final density matrix is given by applying CPTP maps multiple times. Suppose that we repeat a measurement and a feedback operation $N$ times. Let $(m_{1},m_{2},\cdots,m_{N})$ be a sequence of measurement outcomes through this process. After the $n$th measurement, we perform a feedback operation as
$\hat{\rho}_{m_{1},\cdots,m_{n}}^{\prime}=\hat{U}_{m_{n}}\hat{\rho}_{m_{1},\cdots,m_{n}}\hat{U}_{m_{n}}^{\dagger},$ (6)
@@
$\hat{\rho}_{m_{1},\cdots,m_{n}}=\frac{\hat{M}_{m_{n}}\hat{\rho}_{m_{1},\cdots,m_{n-1}}^{\prime}\hat{M}_{m_{n}}^{\dagger}}{p(m_{n}|m_{1},\cdots,m_{n-1})}$ (7)
is the density matrix after the $n$th measurement and
@@
is the probability of an outcome $m_{n}$, conditioned on the past outcomes $m_{1},\cdots,m_{n-1}$. Then, the final density matrix $\hat{\rho}_{\mathrm{f}}^{(N)}$ averaged over all measurement outcomes is given by
@@
Thus, repeated feedback control is described by the repeated application of a CPTP map. Equation (9) suggests an analogy between periodically driven (Floquet) quantum systems [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31] and feedback-controlled quantum systems: a unitary time-evolution operator for one period is repeatedly applied to a state vector in the former, while a CPTP map is repeatedly applied to a density matrix in the latter (see also Table 1).
The above description can be generalized to the cases in which different measurement and feedback operations are performed multiple times during a single cycle. For example, let us perform a second measurement with measurement operators $\{\hat{M}_{m_{2}}^{\prime}\}$ after the first measurement and the subsequent feedback operation. Then, we perform the second feedback control with unitary operators $\{\hat{U}_{m_{1},m_{2}}^{\prime}\}$, which may depend on the outcomes of the first and second measurements. The overall feedback cycle is described by a CPTP map
@@
where $\hat{K}_{m_{1},m_{2}}^{\prime}=U_{m_{1},m_{2}}^{\prime}\hat{M}_{m_{2}}^{\prime}\hat{U}_{m_{1}}\hat{M}_{m_{1}}$ is the Kraus operator. More general cases with nonunitary feedback operations will be discussed in Sec. VIII.
Let $\mathcal{H}$ be a finite-dimensional Hilbert space of the system. Then, the density matrix $\hat{\rho}$ is an element of the space $\mathcal{L}(\mathcal{H})$ of linear operators on $\mathcal{H}$, which is a Hilbert space with dimension $(\dim\mathcal{H})^{2}$. The Hilbert-Schmidt inner product in $\mathcal{L}(\mathcal{H})$ is denoted by $\langle\hat{A},\hat{B}\rangle:=\mathrm{Tr}[\hat{A}^{\dagger}\hat{B}]$. A CPTP map $\mathcal{E}$ [Eq. (5)] is regarded as an operator acting on the operator space $\mathcal{L}(\mathcal{H})$ (i.e., $\mathcal{E}$ is a superoperator). The adjoint superoperator $\mathcal{E}^{\dagger}$ is defined by
@@
where $\xi_{n}\in\mathbb{C}$ denotes an eigenvalue and $\hat{\rho}_{n}^{R}$ ($\hat{\rho}_{n}^{L}$) is the corresponding right (left) eigenoperator which satisfies the biorthogonal condition $\mathrm{Tr}[(\hat{\rho}_{m}^{L})^{\dagger}\hat{\rho}_{n}^{R}]=0$ for $m\neq n$. By using the eigensystem of the CPTP map $\mathcal{E}$, the density matrix after a feedback loop in Eq. (4) is expanded as
@@
An eigensystem of a CPTP map has some general properties (see Fig. 1(a) for a schematic illustration). First, all the eigenvalues $\xi_{n}$ of a CPTP map satisfies $|\xi_{n}|\leq 1$ because of the contraction property of a CPTP map [76]. An eigenmode $\hat{\rho}_{\mathrm{SS}}$ with eigenvalue $+1$ is a steady state of the dynamics in the sense that $\mathcal{E}(\hat{\rho}_{\mathrm{SS}})=\hat{\rho}_{\mathrm{SS}}$. Second, since a CPTP map preserves the trace of the density matrix, that is, $\mathrm{Tr}[\mathcal{E}(\hat{\rho})]=\mathrm{Tr}[\hat{\rho}]$, all the right eigenmodes except for those corresponding to steady states are traceless: $\mathrm{Tr}[\hat{\rho}_{n}^{R}]=0\ (\xi_{n}\neq 1)$. Third, a CPTP map preserves the Hermiticity of an operator since
@@
Therefore, eigenvalues of a CPTP map are real or appear as complex-conjugate pairs.

### v2

@@ I Introduction
Topology plays a key role in classification and characterization of diverse physical phenomena that are robust against local disturbances. In particular, topological phases of matter have been discussed in a wide variety of quantum and classical systems, which are characterized by global geometric structures of eigenstates rather than local order parameters [1, 2, 3, 4, 5, 6]. A quintessential example of topological quantum phenomena is the quantum Hall effect, where chiral transport immune to disorder emerges along the edges of a system [7, 8, 9, 10, 11]. Furthermore, topological phases are enriched by symmetry. For example, time-reversal symmetry permits helical edge states, where time-reversal partners flow in opposite directions with no net mass current, a phenomenon known as the quantum spin Hall effect [12, 13, 14, 15, 16]. Classification of phases of matter through integration of topology and symmetry has been of central interest in condensed matter physics [17, 18, 19].
@@
The general concept of topology can also be utilized to classify nonequilibrium dynamics. Examples include periodically driven (Floquet) systems characterized by topology of unitary time-evolution operators [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31] and topological phases of open systems described by non-Hermitian Hamiltonians [32, 33, 34, 35, 36, 37, 38]. Notably, out-of-equilibrium driving not only provides a useful tool for controlling topological phases, but also induces genuinely dynamical topological phases that have no static counterpart [22, 39, 28, 33]. Finding a new control method for topological phases thus expands the scope of research of topological phases of matter.
@@
Feedback control provides a versatile tool for a wide range of applications in science and technology [40]. In particular, in the research arena of high-precision measurement and control of quantum systems, feedback control has achieved suppression of quantum noise [41], stabilization of desired states [42, 43], cooling [44, 45, 46], and quantum error correction [47, 48, 49, 50]. Feedback control also allows operations that would otherwise be impossible. A prime example is Maxwell’s demon, which was originally conceived as an intelligent being that can decrease entropy of a system against the second law of thermodynamics [51]. In modern parlance, Maxwell’s demon can be formulated as a feedback controller [52, 53], and various types of Maxwell’s demon have experimentally been realized in both classical [54, 55, 56, 57] and quantum [58, 59, 60, 61, 62, 63] systems. In view of versatile applicability of feedback control, we are naturally led to the following question: can we utilize feedback control to create and manipulate nonequilibrium topological phases? This is the problem we wish to address in this paper.
@@
In this paper, we present a class of dynamical topological phases that can be realized with feedback control of quantum systems. Here, a unique role played by feedback control is that one can choose operations depending on measurement outcomes. This flexibility enables versatile real-time control of topological phases that would be hard to achieve by unitary dynamics in which the driving protocol is predetermined and cannot depend on the state of a system under control. Furthermore, feedback control allows active control of a quantum system, in contrast with passive control of open quantum systems by engineered dissipation.
@@
Figure 1: (a) Schematic illustration of the eigenspectrum $\{\xi\}$ (shown by a blue curve and two blue dots) of a quantum channel for topological quantum feedback control. The spectrum involves a closed loop characterized by a winding number and the highly degenerate zero eigenvalues due to local measurement through which nonlocal coherence is projected out. The unit eigenvalue corresponds to a steady state. (b) Discrete quantum feedback control is comprised of a quantum measurement and a subsequent feedback operation that depends on the measurement outcome. A topologically nontrivial winding of the eigenspectrum of a quantum channel [see (a)] implies chiral (unidirectional) transport of a particle on a lattice due to feedback control.
@@
Table 1: Comparison of different classes topological phases. The first row shows five classes of topological phases characterized by the topology of relevant operators shown in the second row. The third row indicates whether the dynamics is static, discrete, or continuous. The fourth row shows whether the system is closed or open. The fifth row shows whether the state of the system is pure or mixed.
@@
Quantum feedback control creates a class of nonequilibrium open quantum systems distinct from those described by non-Hermitian Hamiltonians [33, 34, 35] and Lindblad master equations [64, 65, 66, 67]. We summarize the comparison of different classes of topological phases in Table 1. In particular, discrete quantum feedback control concerns discrete time evolution from an initial quantum state to a post-feedback quantum state. Such time evolution cannot, in general, be expressed by using a generator of continuous time evolution. In fact, as shown in Sec. III, CPTP maps with local Kraus operators have highly degenerate zero eigenvalue, and therefore cannot be described by using any generator [see Eq. (39)]. Physically, the existence of zero eigenvalues can be understood as a consequence of local measurement through which nonlocal coherence in an extended system is projected out [see Fig. 1(a) and Sec. III]. A non-Hermitian Hamiltonian can only describe the dynamics of pure states and thus requires postselection of particular measurement outcomes [68]. The Lindblad quantum master equation can describe the dynamics of mixed states but does not refer to measurement outcomes [40]. In contrast, nonequilibrium topological phases created by quantum feedback control are realized by a nontrivial interplay between quantum measurement and feedback operations. We note that non-Hermitian topology has been applied to quantum channels in Ref. [33]. However, its application beyond zero-dimensional systems is nontrivial and has remained elusive; in fact, the locality of Kraus operators which is relevant in higher dimensions, is the key for defining topological invariants, as discussed in Sec. III.
@@
We present prototypical models of topological feedback control by using quantum Maxwell’s demons, which we shall refer to as topological Maxwell’s demons. In one dimension, a topological Maxwell’s demon causes feedback-assisted chiral transport due to topology. We find that the topological Maxwell’s demon creates topologically distinct phases by changing the duration of feedback control. We also show that feedback control can induce the non-Hermitian skin effect, which is a hallmark of dynamical topology unique to open systems [69, 70, 33, 71, 72, 73, 38, 74]. The chiral transport caused by the topological Maxwell’s demon is protected by topology and hence robust against disorder, noise, and decoherence during the feedback process.
@@
Symmetry plays a pivotal role in classification of topological phases [18, 17, 19, 27, 28, 33, 34, 35, 65, 37, 66, 67]. Here, we perform a symmetry classification of quantum feedback control. We find that only a subset of the 38 symmetry classes of non-Hermitian operators [75, 34, 35] is compatible with quantum channels with projective measurements. We identify ten relevant symmetry classes of quantum channels for discrete quantum feedback control (see Table 2) and provide examples of feedback control for each class. In particular, we propose a quantum Maxwell’s demon with a two-point measurement as an example of symmetry-protected topological feedback control. This demon generates helical spin transport protected by symmetry, which may be regarded as a counterpart of the quantum spin Hall effect in feedback-controlled quantum systems.
@@
Table 2: Ten-fold symmetry classification of discrete quantum feedback control with projective measurement. Each symmetry class is defined according to the presence or absence of $C,K$, and $Q$ symmetries in the CPTP map. Here, $C,K$, and $Q$ symmetries of a CPTP map $\mathcal{E}$ are defined by $\mathcal{V}_{C}\mathcal{E}^{T}\mathcal{V}_{C}^{-1}=\mathcal{E},\mathcal{V}_{K}\mathcal{E}^{*}\mathcal{V}_{K}^{-1}=\mathcal{E}$, and $\mathcal{V}_{Q}\mathcal{E}^{\dagger}\mathcal{V}_{Q}^{-1}=\mathcal{E}$, respectively, where $\mathcal{V}_{X}$ ($X=C,K,Q$) is a unitary superoperator. In the table, 00 indicates the absence of symmetry, and ±$\pm 1$ indicates the presence of symmetry with $\mathcal{V}_{X}\mathcal{V}_{X}^{*}=\pm\mathcal{I}$, where $\mathcal{I}$ denotes the identity superoperator. The labeling of the symmetry classes follows Ref. [34], where psH denotes pseudo-Hermiticity and the subscript $+$ (--) of psH indicates that the unitary superoperator for $C$ or $K$ symmetry commutes (anticommutes) with that of the pseudo-Hermiticity. In addition to the symmetry classes of a CPTP map $\mathcal{E}$, the corresponding Altland-Zirnbauer classes of $i\mathcal{E}$ are shown in the parentheses. In the fifth column, examples of conditions for feedback (FB) control in each symmetry class are shown (see Sec. VI for detailed discussions). CPTP maps with $C$ or $Q$ symmetry can be realized with feedback control with a two-point measurement (2PM). Symmetry classes without $K$ symmetry with $\mathcal{V}_{K}\mathcal{V}_{K}^{*}=+\mathcal{I}$ require unitary symmetry (US) due to the intrinsic modular conjugation symmetry of CPTP maps. AUS- denotes antiunitary symmetry whose square gives $-1$. Some symmetry classes may be realized with an additional feedback control that enforces the symmetry condition (see examples in Sec. VII). The rightmost column shows concrete models of topological feedback control, which are presented in Sec. IV and Sec. VII.
@@
AI $+1$ 00 00 US, 2PM, and reciprocity
AI $+$ psH+ (BDI) $+1$ $+1$ $1$ 2PM and reciprocity
AI (D) 00 $+1$ 00 No symmetry or particle-hole symmetry Chiral Maxwell’s demon / SSH demon
AI $+$ psH- (DIII) $-1$ $+1$ $1$ 2PM, reciprocity, and additional FB Helical Maxwell’s demon
AII $-1$ 00 00 US, 2PM, reciprocity, and additional FB $\mathbb{Z}_{2}$ demon
AII $+$ psH+ (${}^{\dagger})$ $-1$ $-1$ $1$ US, 2PM, AUS-, reciprocity, and additional FB
AII (C) 00 $-1$ 00 US and AUS-
AII $+$ psH- (CI) $+1$ $-1$ $1$ US, 2PM, AUS-, and reciprocity
The rest of this paper is organized as follows. In Sec. II, we introduce quantum channels that describe discrete quantum feedback control. In Sec. III, we use the Bloch theory of translationally invariant quantum channels to show that topological invariants of quantum channels can be defined when Kraus operators are local in space. In Sec. IV, we construct an example of topological feedback control, which shows feedback-induced chiral transport due to its point-gap topology. In Sec. V, we examine symmetry of discrete quantum feedback control and identify ten symmetry classes that are compatible with projective measurement. We also discuss how to construct feedback control in each symmetry class in Sec. VI. On the basis of the symmetry classification, we provide various models of symmetry-protected topological feedback control in Sec. VII. In particular, we present topological feedback control protected by transposition symmetry which exhibits helical spin transport due to an interplay between symmetry and topology. In Sec. VIII, we discuss topological phenomena arising from complete positivity and a trace-preserving property, which are unique to quantum channels and absent in general non-Hermitian systems. In Sec. IX, we extend our theory to the case of nonunitary feedback control, thereby showing robustness of topological feedback control against dissipation. We also propose feedback control with engineered dissipation, where dissipation is used as an essential tool to achieve feedback control. We conclude this paper in Sec. X with an outlook. Some technical details are relegated to the Appendices to avoid digressing from the main subject. In Appendix A, we present the detail of the Bloch theory of quantum channels. In Appendix B, we describe the Bloch theory of quantum feedback control with projective measurement. In Appendix C, we discuss a spectral property of quantum channels for projective measurement. In Appendix D, we summarize the correspondence between symmetry of time-evolution operators and that of their generators. In Appendix E, we provide a supplementary discussion of symmetry-preserving boundary conditions for a $\mathbb{Z}_{2}$ Maxwell’s demon. In Appendix F, we present a model of topological feedback control that shows a non-Hermitian skin effect but with no boundary-localized steady state.
@@ II Discrete quantum feedback control
We consider discrete quantum feedback control in which a feedback cycle consists of quantum measurements and subsequent unitary operations that are conditioned on measurement outcomes. Let $\hat{\rho}_{\mathrm{i}}$ be the initial density matrix of a system. We first perform a quantum measurement described by a set of measurement operators $\{\hat{M}_{m}\}$ subject to the completeness relation
@@
where $\hat{H}_{m}(t)$ is the feedback Hamiltonian which depends on the measurement outcome $m$ and $\mathcal{T}$ is the time-ordering operator. We set $\hbar=1$ throughout this paper. If we take an ensemble average over measurement outcomes, then the final density matrix after one feedback loop is given by [52]
@@
If a feedback cycle is repeated, the final density matrix is given by applying CPTP maps multiple times. Suppose that we repeat a measurement and a feedback operation $N$ times. Let $(m_{1},m_{2},\cdots,m_{N})$ be a sequence of measurement outcomes through this process. The density matrix after the $n$th measurement is given by
$\hat{\rho}_{m_{1},\cdots,m_{n}}=\frac{\hat{M}_{m_{n}}\hat{\rho}_{m_{1},\cdots,m_{n-1}}^{\prime}\hat{M}_{m_{n}}^{\dagger}}{p(m_{n}|m_{1},\cdots,m_{n-1})},$ (6)
@@
is the probability of an outcome $m_{n}$, conditioned on the past outcomes $m_{1},\cdots,m_{n-1}$. After the $n$th measurement, we perform a feedback operation and obtain the postfeedback density matrix as
$\hat{\rho}_{m_{1},\cdots,m_{n}}^{\prime}=\hat{U}_{m_{n}}\hat{\rho}_{m_{1},\cdots,m_{n}}\hat{U}_{m_{n}}^{\dagger}.$ (8)
Then, the final density matrix $\hat{\rho}_{\mathrm{f}}^{(N)}$, which is obtained as the averaged density matrix over all measurement outcomes, is given by
@@
Thus, repeated feedback control is described by repeated applications of a CPTP map. Equation (9) suggests an analogy between periodically driven (Floquet) quantum systems [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31] and feedback-controlled quantum systems: a unitary time-evolution operator for one period is repeatedly acted on a state vector in the former, while a CPTP map is repeatedly acted on a density matrix in the latter (see also Table 1).
The above description can be generalized to the cases in which different measurement and feedback operations are performed multiple times during a single cycle. For example, suppose that we perform a second measurement described by measurement operators $\{\hat{M}_{m_{2}}^{\prime}\}$ after the first measurement and the subsequent feedback operation. Then, we perform the second feedback control with unitary operators $\{\hat{U}_{m_{1},m_{2}}^{\prime}\}$, which may depend on the outcomes of the first and second measurements. The overall feedback cycle is described by a CPTP map
@@
where $\hat{K}_{m_{1},m_{2}}^{\prime}=U_{m_{1},m_{2}}^{\prime}\hat{M}_{m_{2}}^{\prime}\hat{U}_{m_{1}}\hat{M}_{m_{1}}$ is the Kraus operator. More general cases with nonunitary feedback operations will be discussed in Sec. IX.
Let $\mathcal{H}$ be a finite-dimensional Hilbert space of the system. Then, the density matrix $\hat{\rho}$ is an element of the space $\mathcal{L}(\mathcal{H})$ of linear operators on $\mathcal{H}$, which is a Hilbert space with dimension $(\dim\mathcal{H})^{2}$. The Hilbert-Schmidt inner product of two operators $\hat{A}$ and $\hat{B}$ in $\mathcal{L}(\mathcal{H})$ is denoted by $\langle\hat{A},\hat{B}\rangle:=\mathrm{Tr}[\hat{A}^{\dagger}\hat{B}]$. A CPTP map $\mathcal{E}$ [Eq. (5)] is regarded as an operator acting on the operator space $\mathcal{L}(\mathcal{H})$ (i.e., $\mathcal{E}$ is a superoperator). The adjoint superoperator $\mathcal{E}^{\dagger}$ is defined by
@@
where $\xi_{n}\in\mathbb{C}$ denotes an eigenvalue, and $\hat{\rho}_{n}^{R}$ and $\hat{\rho}_{n}^{L}$ are the corresponding right and left eigenoperators which satisfy the biorthogonal relation $\mathrm{Tr}[(\hat{\rho}_{m}^{L})^{\dagger}\hat{\rho}_{n}^{R}]=0$ for $m\neq n$. In terms of the eigensystem of the CPTP map $\mathcal{E}$, the density matrix after a feedback loop in Eq. (4) is expanded as
@@
The eigensystem of a CPTP map has some general properties (see Fig. 1(a) for a schematic illustration). First, all the eigenvalues $\xi_{n}$ of a CPTP map satisfies $|\xi_{n}|\leq 1$ because of the contraction property of a CPTP map [76]. An eigenmode $\hat{\rho}_{\mathrm{SS}}$ with eigenvalue $+1$ is a steady state of the dynamics in the sense that $\mathcal{E}(\hat{\rho}_{\mathrm{SS}})=\hat{\rho}_{\mathrm{SS}}$. Second, since a CPTP map preserves the trace of the density matrix, that is, $\mathrm{Tr}[\mathcal{E}(\hat{\rho})]=\mathrm{Tr}[\hat{\rho}]$, all the right eigenmodes except for those corresponding to steady states are traceless: $\mathrm{Tr}[\hat{\rho}_{n}^{R}]=0\ (\xi_{n}\neq 1)$. Third, a CPTP map preserves the Hermiticity of an operator since
@@
Therefore, eigenvalues of a CPTP map are real or appear in complex-conjugate pairs.